生命テクノサイエンス論文輪講 2 単位 (必修)

Practice for undrstanding scientific papers in bioogical technology  生命テクノサイエンスコース教員

【授業目的】修士課程の研究を行うための論文の読解法を学ぶ．【授業概要】修士論文の研究テーマに関連する論文を読み，所属する研究室にお

いて，発表，討論することを通じて，研究の現状，課題，動向についてまとめる．この課程でプレゼンテーション能力を高めるとともに，今後の研究の展開と社会的意義について理解する．本科目は，工業に関する科目である．

【授業形式】講義および演習【キーワード】論文購読，プレゼンテーション【先行科目】 [先行科目] 【関連科目】『生命テクノサイエンス演習』 (0.5)，『生命テクノサイエンス特別

実験』 (0.5)

【履修要件】なし【履修上の注意】なし【到達目標】論文の内容を解読する方法を学ぶ【授業計画】

1. 発表と議論 1

2. 発表と議論 2

3. 発表と議論 3

4. 発表と議論 4

5. 発表と議論 5

6. 発表と議論 6

7. 発表と議論 7

8. 発表と議論 8

9. 発表と議論 9

10. 発表と議論 10

11. 発表と議論 11

12. 発表と議論 12

13. 発表と議論 13

14. 発表と議論 14

15. 発表と議論 15

16. 発表と議論 16 【成績評価基準】発表の相互評価【教科書】なし【参考書】関連論文【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216747 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 生物事務室(M703) 【備考】各々 1 時間の予習と復習が，各 2 時間の授業の理解と単位取得のために

必要である.

Practice for undrstanding scientific papers in bioogical technology 2 units (compulsory)

Teacher of course

Target⟩ To learn how to read scientific papers for research Outline⟩ In order to obtain information on research and to learn how to present

research results, students read papers related to their subjects of master thesis, present its contents, and discuss about contents of the paper.

Style⟩ Lecture and excercise �������⟩ reading of papar, presentation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Seminar in Biological Technology”(0.5), “Biological science

and technology laboratory”(0.5) Requirement⟩ N/A Notice⟩ N/A Goal⟩ To learn how to read papers

Schedule⟩

1. presentation and discussion 1

2. presentation and discussion 2

3. presentation and discussion 3

4. presentation and discussion 4

5. presentation and discussion 5

6. presentation and discussion 6

7. presentation and discussion 7

8. presentation and discussion 8

9. presentation and discussion 9

10. presentation and discussion 10

11. presentation and discussion 11

12. presentation and discussion 12

13. presentation and discussion 13

14. presentation and discussion 14

15. presentation and discussion 15

16. presentation and discussion 16 Evaluation Criteria⟩ Mutual evaluation of presentation Textbook⟩ N/A Reference⟩ Related papers

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216747

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Administration Office(M703)

Note⟩ Preparation and review of 1 hour each are required for understanding of every class with two hours and taking the credits.